Consider a steam power plant operating on the simple Rankine cycle. Steam enters the turbine

at 16 MPa and 600°C and is condensed in the condenser at a pressure of 8 kPa. Considering an

isentropic efficiency of 85% for both the turbine and pump, determine (a) the net work produced

in the cycle, (b) the heat input in the boiler, (c) the heat rejection in condenser and (d)the thermal

efficiency of this power plant.


1
Expert's answer
2021-07-06T01:15:47-0400

(a) The net work produced in the cycle

Turbine work - Pump work

Turbine work "=h_1-h_2= 3214.37-2455.87= 758.5 kJ\/kg"

Pump work "=h_4-h_3= 421.5-417.43=4.07 kJ\/kg"

Net work = "758.5-4.07 = 754.43 kJ\/kg"

Considering the efficiency "\frac{85}{100}*754.43 = 641.265 kJ/kg"


(b) The heat input in the boiler

"Q_s = h_1-h_4= 3214.37-421.5 =2792.87 kJ\/kg"


(c) The heat rejection in condenser

"Q_R= h_2-h_3=2455.87-417.43=2038.44kJ\/kg"


(d)The thermal efficiency of this power plant.

"\frac{W_{net}}{Q_s}*100=\frac{641.265-4.07}{2792.87}*100=22.82 \%"